OpenStudy (anonymous):

Use a half-angle identity to find the exact value of sin75

OpenStudy (freckles):

hint 150/2=75

OpenStudy (xapproachesinfinity):

\[\sin (75)=\pm \sqrt{\frac{1-\cos150}{2}}\]

OpenStudy (xapproachesinfinity):

that should be clear now!

OpenStudy (xapproachesinfinity):

cos (150)=-cos(30)